OpenStudy (anonymous):

Let A∈ Μ4,3 (that is, a 4 x 3 matrix). Let v1, v2 ∈ R^4 and let w = v1 +v2. Suppose there exists u1, u2 as an element of R^3 such that v1 = Au1 and v2= Au2 prove the Ax=w is consistent

OpenStudy (cruffo):

Question: True of False A matrix equation, Ax= b, is "consistent" if it has at least one solution.

OpenStudy (anonymous):

true

OpenStudy (cruffo):

Ok... we seem to be given alot to work with. We're told that we can assume that there exists a u1 and u2 so that v1 = Au1 and v2 = Au2 Sooo... w = v1 + v2 = Au1 + Au2

OpenStudy (cruffo):

Question: Au1 + Au2 = A(u1 + u2) ???

OpenStudy (anonymous):

that's what i wrote down...

OpenStudy (cruffo):

Right. If Au1 + Au2 = A(u1 + u2), then just take x = u1 + u2 as a solution. To convince yourself that Au1 + Au2 = A(u1 + u2), do an "element-by-element" proof

OpenStudy (anonymous):

that's all i had to do?

OpenStudy (cruffo):

I think so. They may be trying to be tricky with the dimensions. A has to act on a vector in R3, and will return a vector in R4. So the dimensions look fine.

OpenStudy (anonymous):

ok thanks